(Ignore income taxes in this problem.) Allen Company's required rate of return is 14%. The company is considering the purchase of a new machine that will save $10,000 per year in cash operating costs. The machine will cost $40,000 and will have an 8-year useful life with zero salvage value.


Required:


i)     The company would like to use NPV to evaluate the project now. Compute the machine's NPV, assuming cost of capital is 10%. Would you recommend purchase of the machine? Explain.







1
Expert's answer
2021-11-10T10:01:57-0500

Machines NPV:


Discount rate (k)=cost of capital="10\%=0.1"


"\therefore 1+k=1+0.1=1.10"


NPV="-40000+\frac{{10000(1.10^{8} -1)}}{0.10\times1.10^{8}}"


NPV=-40000+53349


="\$13349"


Since NPV is positive, the machine should be purchased